• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 07:57:15 on June15,2002

    Entry number 86075

    keyword=owl shift summary

    Shift crew: R. Michaels, G. Batigne, H. Ibrahim, S. Frullani

    Start shift doing HARP scans, CODA runs 1492 - 1494

    Start kinematics Q2D_c20 with
    P_L = 3.0267 theta_L = 22.148
    P_R = 2.4144 theta_R = 31.420
    E_beam = 4.703 (Tiefenback)

    CODA run........Comments

    1495.................Junk. At the same time:

    IA and PZT feedback tests commence (R. Suleiman, R. Michaels)


    1496.................Production Run.... 1 M events
    1497.................Production.....600 K events
    1498.................Production.....400 K events
    1499.................Efficiency study of R-arm (ps1 = ps2 = 1)

    Next, we change to kinematics Q2D_c40
    R-arm changed to
    P_R = 2.3493 theta_R = 40.48

    1500.................LD2 target.... 58 K events......Good run, but:
    Oops, we forgot to run trigsetup (no harm since time difference between
    spectrometers is, and was, smaller than least count.). Run trigsetup now.
    1501.................Optics Target
    Finish IA/PZT feedback tests. As expected, they did no harm.
    1502.................LD2 production....630 K....Good, but ended with CODA crash.
    1503.................Production Run.....1 M
    Beam current increased to 65 uA.
    1504.................Production Run.....1 M
    1505.................Production Run.....1 M
    1506.................Production Run......in progress

    Doing checklist we noted some presumably minor problems:
    minor problems in checklist